# Teams of LLM Agents can Exploit Zero-Day Vulnerabilities

### v1

@@ 9 Conclusions and Limitations
In this work, we show that teams of LLM agents can autonomously exploit zero-day vulnerabilities, resolving an open question posed by prior work [5]. Our findings suggest that cybersecurity, on both the offensive and defensive side, will increase in pace. Now, black-hat actors can use AI agents to hack websites. On the other hand, penetration testers can use AI agents to aid in more frequent penetration testing. It is unclear whether AI agents will aid cybersecurity offense or defense more and we hope that future work addresses this question.
Beyond the immediate impact of our work, we hope that our work inspires frontier LLM providers to think carefully about their deployments.
@@
Acknowledgements
We would like to acknowledge the Open Philanthropy project for funding this research in part.

### v2

@@ 9 Conclusions
In this work, we show that teams of LLM agents can autonomously exploit zero-day vulnerabilities, resolving an open question posed by prior work Fang et al. (2024a). Our findings suggest that cybersecurity, on both the offensive and defensive side, will increase in pace. Now, black-hat actors can use AI agents to hack websites. On the other hand, penetration testers can use AI agents to aid in more frequent penetration testing. It is unclear whether AI agents will aid cybersecurity offense or defense more and we hope that future work addresses this question. Beyond the immediate impact of our work, we hope that our work inspires frontier LLM providers to think carefully about their deployments.
10 Limitations, Ethical Considerations
@@
A major consideration when conducting research in potentially harmful uses of LLMs is that malicious actors can use the ideas for nefarious purposes. To help alleviate such issues, we have elected not to release our code or prompts publicly as OpenAI has requested that we keep our agents confidential. This is in line with prior work Fang et al. (2024b, a) and best practice for cybersecurity OWASP (2024). Furthermore, we have disclosed our findings to OpenAI as part of their responsible disclosure program.